simplifly 1/3-(2/5+1/4)/(3/8+1/3)=
Answer by tommyt3rd(5050)

1/3-(2/5+1/4)/(3/8+1/3)=
1/3-(8/20+5/20)/(9/24+8/24)=
1/3-(13/20)(17/24)=-149/255